I need help with Age and Quotient- Remainder word problems.
Here is my question:
A girl is now one-third as old as her mother, and in three years she will be two-fifths as old as her mother is then. Find their present ages.

Found 2 solutions by stanbon, rothauserc:
Answer by stanbon(75887)

A girl is now one-third as old as her mother, and in three years she will be two-fifths as old as her mother is then. Find their present ages.
-----------
Equations:
g = (1/3)m
g+3 = (2/5)(m+3)
----------------------
Substitute for "g" and solve for "m"::
(1/3)m + 3 = (2/5)m + 6/5
-------
Multiply thru by 15 to get rid of the denominators:
5m + 45 = 6m + 18
m = 27 (mom's age now)
-----
Solve for "g"::
g = (1/3)m
g = (1/3)27 = 9 (girl's age now)
----------------
Cheers,
Stan H.
-------------

Answer by rothauserc(4718)

we have two equations in two unknowns, let x be the girl and y be her mother, then
x = y/3
x + 3 = (2/5) * (y + 3)
substitute for x in second equation
y/3 + 3 = (2/5) * (y + 3)
y/3 + 3 = 2y/5 + 6/5
multiply both sides of = by 15
5y + 45 = 6y + 18
y = 27 and x = 9
the girl is 9 and her mother is 27
check the answer by substituting for x and y in the second equation
9 + 3 = 2/5 * (27 + 3)
12 = (2/5) * 30
12 = 12
answer checks